Find the equation of the line having the given slope and containing the given points. Express your answer in the form x=a, y=b, or y=mx+b.
m=-5,(5,0)
The equation of the line is y=?
Help me please.

Answer by scott8148(6628)

the point-slope form of a line with slope m through point (c,d) is y-d=m(x-c)

y-0=-5(x-5) __ y=-5x+25